order from least to greatest
-8/9 , -7/8 , -22/25

Answer by checkley75(3666)

-8/9=-.888888888 least
-7/8=-.875 greatest
-22/25=.880 middle
-.88888888,-.880,-.875
------------------------------------------------------
-8/9=1600/1800
-7/8=1575/1800
-22/15=1584/1800
-1600/1800,-1584/1800,-1575/1800